OpenStudy (anonymous):

Solve by the substitution method. What is the ordered pair? 6m + n = 34 and m - 5n = 47.

OpenStudy (anonymous):

n=34-6m. m-170+30m=47, 31m=217, m=7. n=-8

myininaya (myininaya):

if m-5n=47, then m=47+5n so we have 6(47+5n)+n=34 282+30n+n=34 31n=34-282 31n= -248 n=-248/31 n=-8 so m=47+5n=47+5(-8)=7

OpenStudy (anonymous):

So would the ordered pair be (-8, 7)? Or (7,-8)?